OpenStudy (kisforkayla):

Solve the following system of equations: −2x + y = 1 −4x + y = −1 (3, 1) (−1, 3) (−1, −3) (1, 3)

MsBrains (ms-brains):

Any ideas? :)

OpenStudy (wolf1728):

a) −2x + y = 1 b) −4x + y = −1 Multiply equation "a" by -2 then add the equations a) 4x - 2y = -2 b) −4x + y = −1 -y = -3 y = 3

OpenStudy (kisforkayla):

thanks

OpenStudy (wolf1728):

u r welcome